OpenStudy (albert0898):

34. Find four consecutive odd integers such that five times the sum of the first and third integers exceeds four times the sum of the second and last integers by 14. Explain.

OpenStudy (anonymous):

treat the first integer as x, the next as x + 2, and the third as x + 4

OpenStudy (p0sitr0n):

x x +2 x+4 whre x is odd 5(x + x+4) = 4(x+2 +x+4) +14

OpenStudy (p0sitr0n):

10x +20 = 8x + 12 +14 2x = 6 x= 3 5 7

OpenStudy (anonymous):

@P0sitr0n should have a '>' sign?

OpenStudy (p0sitr0n):

read till the end.

OpenStudy (anonymous):

i read the word 'exceeds'

OpenStudy (p0sitr0n):

by 14, so its 14 plus?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3, 5, and 7 do not seem to work?

OpenStudy (albert0898):

18, 20, 22 is the answer, but how...

OpenStudy (anonymous):

these are not odd integers...?

OpenStudy (p0sitr0n):

lol true

OpenStudy (anonymous):

@P0sitr0n, seems to be an arithmetic error above

OpenStudy (anonymous):

5(x + x+4) = 4(x+2 +x+4) +14 does not equal 10x +20 = 8x + 12 +14

OpenStudy (albert0898):

Oh wait sorry, the answer is 13, 15, 17, 19

OpenStudy (albert0898):

And it's 4 consecutive odd integers.

OpenStudy (anonymous):

i get 11, 13, 15

OpenStudy (anonymous):

the right hand value of the above equation should be 10x +20 = 8x + 24 +14

OpenStudy (anonymous):

solve that and x will equal 9, but since you need to exceed, you can use 9, you have to move up one odd integer.

OpenStudy (anonymous):

you CAN'T use 9, sorry

OpenStudy (albert0898):

Alright, can you list the steps from the beginning?

OpenStudy (anonymous):

@Albert0898, do you see how we arrived at the equation?

OpenStudy (albert0898):

No, this one is kind of confusing to me.

OpenStudy (anonymous):

okay, I will try to explain. you need three consecutive odd integers. start with the first one, we called that x. the next would be 'x + 2', and the the last one would be 'x + 4' still with me?

OpenStudy (albert0898):

Yes

OpenStudy (anonymous):

okay, then the problem calls for the FIVE TIMES THE SUM of the first and third numbers, which are x and 'x + 4' to EXCEED FOUR TIMES THE SUM of the second and last integers ('x + 2' and 'x + 4') by 14.

OpenStudy (anonymous):

we arrive at this inequality:\[5(x + x + 4) > 4(x + 2 + x + 4) + 14\]

OpenStudy (anonymous):

\[5x + 5x + 20 > (4x + 8 + 4x + 16) + 14\]

OpenStudy (albert0898):

AHHHHH, I see now!! I wasn't sure if I was allowed to put a greater than sign because I couldn't figure out how I'd get the equivalent to 'x'. But please do continue

OpenStudy (anonymous):

\[10x + 20 > 8x + 38\]\[2x > 18\]\[x > 9\]

OpenStudy (anonymous):

glad to know you see it. wonderful! since x has to be GREATER than 9, we have to move to 11

OpenStudy (albert0898):

So x has to be greater than 9? There's an infinite amount of choices for that.

OpenStudy (anonymous):

is there?

OpenStudy (albert0898):

The answer is 13, 15, 17, 19 ~ I sort of get it now

OpenStudy (anonymous):

good. i guess there are an infinite amount. you only need three integers and 11 is the first you can start with

OpenStudy (anonymous):

and thanks for the medal. good luck

OpenStudy (albert0898):

You're Welcome and Thank You!